The first four moments of a distribution about the value 4 of the variable are -1.5,17,-30 and 108. Find the moments about mean, 1 β and 2 β . Find also the moments about (i) the origin, and (ii) the point x=2.
1
Expert's answer
2020-03-11T11:41:18-0400

"i) \\beta_n=M(\\xi-M\\xi)\\text{ --- the n-th central moment}.\\\\\n1) M(\\xi-4)=-1.5\\\\\nM\\xi=4-1.5=2.5\\\\\nM(\\xi-M\\xi)=0\\text{ --- the first cental moment}.\\\\\n2) M(\\xi-4)^2=17\\\\\nM(\\xi-M\\xi)^2=M\\xi^2-(M\\xi)^2\\\\\nM\\xi^2-4^2=17\\\\\nM\\xi^2=33\\\\\nM(\\xi-M\\xi)^2=33-(2.5)^2=26.75\\text{ --- the second central moment}.\\\\\n3) M(\\xi-4)^3=-30\\\\\nM(\\xi-M\\xi)^3=M(\\xi^3-3\\xi^2M\\xi+3\\xi(M\\xi)^2-(M\\xi)^3)=M\\xi^3-3M\\xi^2M\\xi+3(M\\xi)^3-(M\\xi)^3=M\\xi^3-3M\\xi^2M\\xi+2(M\\xi)^3.\\\\\nM(\\xi-4)^3=M\\xi^3-3\\cdot 33\\cdot 4+2\\cdot 4^3=-30\\\\\nM\\xi^3=238.\\\\\nM(\\xi-M\\xi)^3=238-3\\cdot 33\\cdot 2.5+2\\cdot (2.5)^3=21.75\\text{ --- the third central moment}.\\\\\n4) M(\\xi-4)^4=108.\\\\\nM(\\xi-M\\xi)^4=M(\\xi^4-4\\xi^3M\\xi+6\\xi^2(M\\xi)^2-4\\xi(M\\xi)^3+(M\\xi)^4)=\\\\\n=M\\xi^4-4M\\xi^3M\\xi+6M\\xi^2(M\\xi)^2-4(M\\xi)^4+(M\\xi)^4=\\\\\n=M\\xi^4-4M\\xi^3M\\xi+6M\\xi^2(M\\xi)^2-3(M\\xi)^4.\\\\\nM(\\xi-4)^4=M\\xi^4-4\\cdot 238\\cdot 4+6\\cdot 33\\cdot 4^2-3\\cdot 4^4=108.\\\\\nM\\xi^4=1516.\\\\\nM(\\xi-2.5)^4=1516-4\\cdot 238\\cdot 2.5+6\\cdot 33\\cdot (2.5)^2-3\\cdot(2.5)^4=256.3125\\text{ --- the fourth central moment}.\\\\\n\\sigma\\approx 5.17.\\\\\nA=\\frac{M(\u03be\u2212M\u03be)^3}{\\sigma^3}\u200b=\\frac{21.75}{138.35}=0.157 \u2014 \\text{ asymmetry of the distribution}.\\\\\nE=\\frac{M(\u03be\u2212M\u03be)^4\u200b}{\\sigma^4}=\\frac{256.3125}{715.5625}\u200b=0.358 \u2014 \\text{ kurtosis of the distribution}.\\\\\n\\text{Moments about the origin:}\\\\\nM\\xi=2.5\\\\\nM\\xi^2=33\\\\\nM\\xi^3=238\\\\\nM\\xi^4=1516."